OpenStudy (anonymous):

K is the midpoint of . The coordinates of K are (3, 5) and the coordinates of L are (-5, -6), find the coordinates of M. quick medal

pooja195 (pooja195):

use the midpoint formula \[\huge~\rm~\frac{ x_2+x_1 }{ 2},\frac{ y_2+y_1 }{ 2 } \]

pooja195 (pooja195):

\[\huge~\rm~\frac{ -5+3 }{ 2},\frac{ -6+5 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ -5+3 }{ 2 } = -1\]

pooja195 (pooja195):

Not quite... We want coordinates -5+3=?

OpenStudy (anonymous):

\[\frac{ -6+5 }{ 2 } = 0.5\]

OpenStudy (anonymous):

yes

pooja195 (pooja195):

oh ok i see what you did

OpenStudy (anonymous):

( -1, 0.5 )

pooja195 (pooja195):

Correct you can leave it as (-1,-1/2) if you wish it doesnt matter

OpenStudy (anonymous):

thanks